the difference between two numbers is 10.When eight times the smaller number is added to 40, the result i ten less than three times the larger number. find the numbers
Found 2 solutions by nyc_function, lwsshak3:
Answer by nyc_function(2741)

Let x and y be the numbers. You can use any letters.
Let x = larger number
Let y = smaller number
x - y = 10
8x + 40 = 3y - 10

You have a system if linear equations in two variables.

I will solve x - y = 10 for x and plug the answer into the other equation to find y.

x - y = 10
x = y + 10
8(y + 10) + 40 = 3y - 10
8y + 80 + 40 = 3y - 10
8y + 120 = 3y - 10
8y - 3y = -120 - 10

5y = -130

y = -130/5

y = -26

To solve for x, replace y with -26 in either of the two equations above and simplify.

I will use x - y = 10.

Let y = -26

x - (-26) = 10

x + 26 = 10
x = 10 - 26

x = -16

The smaller number is -26 and the bigger number is -16.

The number -16 is BIGGER than -26 because it is closer to zero on the real number line.




Answer by lwsshak3(11628)

the difference between two numbers is 10.When eight times the smaller number is added to 40, the result i ten less than three times the larger number. find the numbers
**
x=smaller no.
y=larger no.
x-y=10
y=x+10
..
8x+40=3y=3(x+10)-10=3x+30-10
5x=-50+30=-20
x=-4
y=x+10=6
..
ans:
smaller no.=-4
larger no.=6